OpenStudy (anonymous):

How would I do this one? From a balloon 760 feet high, the angle of depression to the ranger headquarters is 89°18'. How far is the headquarters from a point on the ground directly below the balloon (to the nearest foot)?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What answer did you get?

OpenStudy (anonymous):

These were the choices.. -1 ft 4 ft 7 ft 9 ft But I don't know how to work this one. It would be wanted in degrees.. Right?

jimthompson5910 (jim_thompson5910):

Hint: see this attached image

OpenStudy (anonymous):

That's what I thought.. But you don't do tan(89.3).. THat's normally how I do them, so I'm lost.

jimthompson5910 (jim_thompson5910):

89°18' = 89 degrees + 18/60 degrees 89°18' = (89 + 18/60) degrees 89°18' = 89.3 degrees so you are correct there

jimthompson5910 (jim_thompson5910):

tan(angle) = opp/adj tan(89.3) = 760/x solve for x

OpenStudy (anonymous):

Oh, okay. Hang on.

OpenStudy (anonymous):

I got ~.108. What the poop?

jimthompson5910 (jim_thompson5910):

tan(89.3) = 760/x x*tan(89.3) = 760 x = 760/tan(89.3) x = ??

OpenStudy (anonymous):

Oops, I'm an idiot. Lol. I did it backwards. I got 9. C:

jimthompson5910 (jim_thompson5910):

yep, 9.28561362328544 which rounds to 9

OpenStudy (anonymous):

I have one more question that is kind of like this. Do you want me to post it seperately?

jimthompson5910 (jim_thompson5910):

Either way works

OpenStudy (anonymous):

A person is watching a boat from the top of a lighthouse. The boat is approaching the lighthouse directly. When first noticed the angle of depression to the boat is 14°52'. When the boat stops, the angle of depression is 45°10'. The lighthouse is 200 feet tall. How far did the boat travel from when it was first noticed until it stopped? Round your answer to the hundredths place.

OpenStudy (anonymous):

I got 554.88. :)

jimthompson5910 (jim_thompson5910):

That's very close, but I'm getting something slightly smaller.

OpenStudy (anonymous):

How did you do it? I did it the way we did the last problem, and then subtracted my answers. I think my problem strats when I try to convert 14(degrees)52' to all degrees. I got 14.86.

jimthompson5910 (jim_thompson5910):

Keep it like 14 + 52/60 I typed this into my calculator 200/tan(14+52/60) - 200/tan(45+10/60)

OpenStudy (anonymous):

That time I got 554.58.

jimthompson5910 (jim_thompson5910):

Me too 200/tan(14+52/60) - 200/tan(45+10/60) = 554.579127143252 which rounds to 554.58

OpenStudy (anonymous):

Yay! Thank you so much! God bless you! c:

jimthompson5910 (jim_thompson5910):

you're welcome